# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v2

@@ Appendix A Derivative Analysis for the Inverted Hinge Loss Function
Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IH}}$. Firstly, for convenience, we define $p_{t}=p_{\theta}(x_{t}|x_{<t})$ and $\hat{p}_{t}=\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$. The loss function can be rewritten as:
$\mathcal{L}_{\text{IH}}({\bm{x}})=\max\left(0,1+p_{t}-\hat{p}_{t}\right)$
To calculate the derivative of $\mathcal{L}_{\text{IH}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$. In the case where $01+p_{t}-\hat{p}_{t}>$ 0, using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ is as follows:
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $1+p_{t}-\hat{p}_{t}\leq 0$. In summary, the derivatives of the loss function $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
$\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$

### v3

@@ Appendix A Derivative Analysis for the Inverted Hinge Loss Function
Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IHL}}$. Firstly, for convenience, we define $p_{t}=p_{\theta}(x_{t}|x_{<t})$ and $\hat{p}_{t}=\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$. The loss function can be rewritten as:
$\mathcal{L}_{\text{IHL}}({\bm{x}})=1+p_{t}-\hat{p}_{t}$
To calculate the derivative of $\mathcal{L}_{\text{IHL}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$. Using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IHL}}$ with respect to $y_{t}^{(v)}$ is as follows:
$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
@@
In summary, the derivatives of the loss function $\mathcal{L}_{\text{IHL}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$